OpenStudy (jh99):

x^2+4y^2=4 y=x+1 Part I: What two shapes intersect in this problem? (2 points) Part II: Use substitution to solve for the two values of x. (4 points) Part III: Use the values of x you just found to get the corresponding values of y, representing the points of intersection of these two shapes. Write your answers in the form (x,y). (4 points)

OpenStudy (photon336):

what you can probably do is this: \[y = (x+1)\] you can plug in x+1 wherever you see y in the first equation. \[x^{2}+4(x+1)^{2} = 4\] \[x^{2}+4(x^{2}+2x+1) = 4\]

OpenStudy (photon336):

\[(5x^{2}+8x+4) = 4\] \[5x^{2}+8x = 0\] do you know how you would continue?

OpenStudy (photon336):

@jh99

OpenStudy (photon336):

How did you solve it? \[x(5x+8) = 0 \]

OpenStudy (jh99):

@photon336 I apologize for being late. Thank you for helping me out. On the equation above, you solved for x?

OpenStudy (photon336):

x(5x+8) = 0

OpenStudy (photon336):

ideas?

OpenStudy (jh99):

Can we say 5x+8=0 ?

OpenStudy (photon336):

yes we can

OpenStudy (photon336):

okay so step #1 solve for x

OpenStudy (jh99):

x=-5/8?

OpenStudy (photon336):

\[5x+8 = 0\] \[\frac{ -8 }{ 5 } = x \]

OpenStudy (jh99):

We need to plug in that value to solve for y?

OpenStudy (jh99):

Also, I was wondering if the first equation is a circle..?

OpenStudy (photon336):

okay so next step let's plug this into the original equation y = x+1

OpenStudy (photon336):

yep :) it is

OpenStudy (jh99):

We'll get the same results if we put it into the first equation, right? (Just to make sure)

OpenStudy (photon336):

\[(y = x+1) = \frac{ -8 }{ 5 }+1 = \frac{ -8+5 }{ 5 } = \frac{ -3 }{ 5 } = y \]

OpenStudy (photon336):

so we've got x = -8/5 and y = -3/5 now to see if this is true let's plug it into the circle equation

OpenStudy (jh99):

Do we need to plug it into a specific equation? It has to be the first one? The second one would be much easier!

OpenStudy (photon336):

so yeah we need to just do this to check. if we plug in x and y into the circle and get 4 then our equation is correct. \[x^{2}+4y^{2} = 4\] \[(\frac{ -8 }{ 5 })^{2}+4(\frac{ -3 }{ 5 })^{2} = 4\] \[\frac{ 64 }{ 25 }+4*(\frac{ 9 }{ 25 }) = \frac{ 36 }{ 25 }+\frac{ 64 }{ 25 } = \frac{ 100 }{ 25 } = 4\]

OpenStudy (jh99):

Thank you so much for helping me out! We can also just plug that in for y=x+1, right ? again, thanks a lot.

OpenStudy (photon336):

wait before you go there is another point we need

OpenStudy (photon336):

note that (-8/5,-3/5) is only one point of intersection between the line and the circle.

OpenStudy (photon336):

I also want to show you how the graph looks and the intersection.

OpenStudy (photon336):

take a look http://prntscr.com/biyjsb

OpenStudy (photon336):

now to find the second point. notice how in the first equation we had: \[(5x^{2}+8x) = x(5x+8) ~ x = 0 \] x = 0 is another solution so we should plug into both equations: \[(y+1) => (0+1) = 1 \] and in \[(x^{2})+(4y^{2}) = 4\] \[(0)^{2}+ 4(1)^{2} = 0+4 = 4 \]

OpenStudy (photon336):

Whenever you solve a system of equations if you find one set of points say x = 0 it must satisfy both equations if it doesn't then your answer is wrong.

Parth (parthkohli):

For part I, you have to know that the first is an ellipse and the second is a line.

OpenStudy (photon336):

yeah also this is a circle in case you're interested. \[x^{2}+4y^{2} = 4 | (x-h)^{2}+(y-k)^{2} = r^{2}| center~is~(h,k)|(0,0)\]

OpenStudy (jh99):

thank you so much guys!

OpenStudy (jh99):

@Photon336 wait, how can a circle be an ellipse?